# Latent space configuration for improved generalization in supervised autoencoder neural networks

### v2

@@ 2.3 Supervised Autoencoders
Whereas AEs were initially proposed for unsupervised learning scenarios, they were proven useful for semi-supervised and supervised learning tasks, too. As any AE, SAE has encoder-decoder architecture with its decoder substituted by a classifier neural network. Decoder and classifier can also be used in parallel so that decoder output is used for regularization [14], but this use case is outside of the scope of this study. For supervised learning the reconstruction loss is substituted with classification loss most commonly being Cross Entropy (CE) loss. In this paper we study LS of SAE with architecture shown in Table 1 where Double layer means two subsequent sequences of convolution, batch normalization [15], and ReLU; and Down layer consists of MaxPool [16] followed by Double layer. This encoder architecture is inspired by VAE architecture [17] without (1) and (2) so dense layer outputs have no statistical meaning and can be referred to as a and b, so LS representation is z = a+b. Classifier is a simple 3-layer fully-connected network. Inference of such SAE is deterministic due to the absence of N(0,1) term in z.
@@ 5.2 Conventional training and LS configuration training
[…]
@@ 6.1 Similarity in LS
$v_{ji}=\begin{cases}1-k_{b}\cdot sin(\frac{d_{ji}}{r_{ci}})\ if\ d_{ji}\leq r_{ci},\\ ReLU(\frac{b_{c}\cdot(R_{d}-d_{ji})}{R_{d}-r_{ci}})\ otherwise;\\ \end{cases}$ (12)

### v3

@@ 2.3 Supervised Autoencoders
Whereas AEs were initially proposed for unsupervised learning scenarios, they were proven useful for semi-supervised and supervised learning tasks, too. As any AE, SAE has encoder-decoder architecture with its decoder substituted by a classifier neural network. Decoder and classifier can also be used in parallel so that decoder output is used for regularization [14], but this use case is outside of the scope of this study. For supervised learning the reconstruction loss is substituted with classification loss most commonly being Cross Entropy (CE) loss. In this paper we study LS of SAE with architecture shown in Table 1 where Double layer means two subsequent sequences of convolution, batch normalization [15], and ReLU; and Down layer consists of MaxPool [16] followed by Double layer. This encoder architecture is inspired by VAE architecture [17] without (1) and (2) so dense layer outputs have no statistical meaning and can be referred to as a and b, so LS representation is z = a+exp(b). Classifier is a simple 3-layer fully-connected network. Inference of such SAE is deterministic due to the absence of N(0,1) term in z.
@@ 5.2 Conventional training and LS configuration training
Geometric loss configurations has the following parameters: five clusters nc = 5 that are at dc = 0.85 distance from center (0,0) and $2\pi/\textit{n\textsubscript{c}}$ = 72° angles relative to the neighbors with cluster radius rc = 0.34. Figures 7 - 9 show positions of LS embeddings of train data (small points) with Figures 8 and 9 also showing Market generalization data LS embeddings (large points). Figure 7 (a) - (c) shows that LCE + LG training method allows to achieve the desired LS distributions for all experiments. Clusters resemble petals due to large number of points and the interplay between LCE and LG. Figure 7 shows that in all cases clusters appear in the desired positions with all training points being roughly inside a circle with radius two. All LS encodings in the Market transfer experiments also lie withing the specified regions not leaving the configured LS. This illustrates that even that LS condition is present only during training, the learned weights succeed in projecting new unseen data into the specified areas in LS.
@@ 6.1 Similarity in LS
[…]